In this question, you will need to find the general solution to the equation [i 25 0 2 3 8 -1 || *2 (1 4 9 1] | x4 in vectorThe general solution is X= y + xw, where V, W are vectors, and X is a variable (X1, X2, X3, or X4). Write the top entry of thWrite the bottom entry of the vector V. Your answer should be an integer. Do not use decimal points. (An answer of the form 1Write the bottom entry of the vector W. Your answer should be an integer. Do not use decimal points. (An answer of the form 1Consider a Markov chain with initial state vector and migration matrix [1 0 P= 1 1What is the long-term trend vector? Write the top entry. Your answer must be a reduced fraction using a slash (e.g., if yourSuppose A is an 3 x 4 matrix such that Ax=b is always) consistent for all vectors be R3. For each of the following, write T f
Answer #1

According to HomeworkLib policy we can solve only one question.

riotol 27 1 x = 9, +%3=2, 12 +223 = -2 4 = 1 4=2-23; 2 = 2-2 na ay 261 2-23 1 . 26 CV- 2- u first entry a first entry of w= a